ZEKE SPECTROSCOPY OF GROUP VIB M(C_{6}H_{6})_{2} (M = Cr, Mo, W) SANDWICH COMPLEXES

\underline{BRADFORD R. SOHNLEIN} and DONG-SHENG YANG, Department of Chemistry, University of Kentucky, Lexington, KY 40506-0055.

Dibenzene complexes of chromium, molybdenum, and tungsten, M(C_{6}H_{6})_{2} (M = Cr, Mo, W), were produced by laser ablation in a supersonic jet and probed by pulsed field ionization-zero electron kinetic energy (PFI-ZEKE) electron spectroscopy. The adiabatic ionization potentials were measured to be 44081(6), 44581(11), and 43635(6) cm^{−1} for Cr(C_{6}H_{6})_{2}, Mo(C_{6}H_{6})_{2}, and W(C_{6}H_{6})_{2}, respectively. The chromium-dibenzene spectrum displays three quanta excitations of the 265 cm^{−1} metal-benzene symmetric stretch mode in the cation, whereas the molybdenum and tungsten complexes show only one excitation of the metal-ligand stretch of 277 and 370 cm^{−1}, respectively. The shift in the ionization energy from the bare metal atom to the complex (10494, 12623 and 19792 cm^{−1}) and metal-ligand stretch frequencies increase, as the group is descended. These observations indicate a stronger metal-benzene interaction for the heavier atoms, especially tungsten. The ground states of the neutral and ionic complexes appear to be in D_{6h} symmetry with the hydrogen atoms slightly bent towards the metal center, as suggested by the good agreement between experiment and theory.